OpenStudy (anonymous):

I need help with this question what is -7.4(10+a)?

OpenStudy (mathmale):

Apply the distributive property of multiplication. Multiply 10 by (-7.4). then multiply "a" by (-7.4). Combine these two results to obtain the final answer.

TheSmartOne (thesmartone):

Hint: a(b+c) = ab + ac

OpenStudy (mathmale):

That's the distributive property of multiplication. Example: 2(x+2y)=2x + 2*2y=2x+4y

OpenStudy (mathmale):

Queenfood: I'd suggest you go ahead and try something. You'd likely learn from doing so, and surely learn more from making mistakes than from remaining silent.

OpenStudy (anonymous):

Ok I just was told that the problem is a expression not a equation but thanx for your guys's help anyway. You learn something no everyday! :)